Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group November 3, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 3, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/03/98 (Tuesday)

Segment 1 (Emidio to ANF): Report for 11/3/98: The 8 mile segment of pipe from EMP's 33.34 to 25.48 was de-watered in order to correct the vapor lock situation. On Tejon Bench, welding and tie-in was underway at EMP 18.48. Shading occurred between 17.53 to 17.46. Welding, x-ray and coating took place from EMP's 17.25 to 17.21. Trench has been excavated to 17.12. Potholing took place at 15.45. Grapevine Road was marked for sawing from EMP's 16.43 to 15.72. Sawing took place from 15.40 to 15.35. At Grapevine Pump Station, pipe installation took place inside the pipe vault. Foundation work continued for the substation. Backfill over the conduit connected from the pipe vault to the electrical building began. Work continued at the I-5/Grapevine bore (EMP 14.62), the California Aqueduct Bore (EMP 12.28) and the Stream Crossing #4 bore (EMP 9.03). On Sabodon Road, backfill occurred at the segment tied-in from 7.06 to 7.19. On Road B long segments of pipe already welded, x-rayed and coated were lowered-in from EMP's 7.04 to 7.60. Pipe was welded, x-rayed and coated from EMP 6.70 to 6.54. Pipe was welded and x-rayed to EMP 6.38. On Hwy. 166, drilling of the Hwy. 166 bore at EMP 1.18 began. Conduit has been placed to EMP .97 and trench has been excavated to .65. Electrical work and pipe fitting continued.

Segment 2 (ANF to Sylmar): 2C (Templin Hwy to Castaic): Work continued on the block wall on Whitaker Reduction Station access road. At the Station, electrical clean-up occurred. Revegetation with curlex was completed at the first gully north of I-5 Bore, EMP 53.41. Revegetation work continued north of EMP 54.65 and south of 54.74. Welding in the trench reached north to EMP 55.69. Tents were used in the afternoon as winds picked up speed. Backfill took place near 56.06. 2B (Castaic to Sylmar): F/O cable was fused at the intersection of Hwy 126 and The Old Road. Surveying took place at Santa Clara River in preparation for the F/O directional drill. F/O pulling assist pits were backfilled at PGA Ranch. At Calgrove Blvd and The Old Road, concrete-coated pipe was inserted into casing, and an additional section of pipe welded on and x-rayed. Backfill material continued to be sorted at Weldon Cyn. Backfill, base rock, asphalt, and F/O crews worked between Sierra Hwy and Sunshine Cyn. Bending and lowering-in took place south of Sunshine Cyn. At Sunshine Cyn, repairs were made on the riser. At Balboa Blvd, asphalt was applied around the vault.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 3 November 1998: 3A: Minor work on the tie in at Verdugo Wash continued today. The other crews worked on San Fernando and Hollywood Way Bridge crossings. 3B: Backfill continued on north Downey Bridge, with razor wire being hung on pipe. 3C: Excavation crews worked late at the Hubbard Avenue bore site to place shields and secure area with fencing. Backfill continued near Grapevine Creek. Ditch crew worked near the tie in to the bore under the railroad spur. Welding, x-ray, and coating progressed with their work on the pipe north of Sepulveda. Some asphalt was removed from San Fernando Road over pipe alignment. The crew at the aqueduct crossing worked on moving forms to next sections to be poured.

Segment 4B (El Segundo to Lynwood) 11-3-98: Backfilling activities continued at 108 Th St. ESEMP 1.16, Santa Ana ESEMP 0.04, Alameda and Fernwood AV ESEMP 0.0. Coating and slurry backfilling took place at Western ESEMP 5.56. Construction activities continued at El Segundo station ESEMP 11.00. Construction activities continued at Lynwood station EMP 0.00.

Segment 5 (Lynwood to Long Beach) Report for 3 November 1998: Directional drill at Highway 91/Compton Creek progressed without problem. Water and mud is being contained on the ROW. Activities in preparation for the bore at Wilmington Refinery continued.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page